Five years ago a father was thrice as old as his son. Fifteen years hence he will be twice as old as his son find the present age of his son (in years)
Answer by ankor@dixie-net.com(22740)

let f = father's present age
let s = son's age
:
write an equation for each statement, simplify as much as possible.
:
Five years ago a father was thrice as old as his son.
f - 5 = 3(s-5)
f - 5 = 3s - 15
f = 3s - 15 + 5
f = 3s - 10
:
Fifteen years hence he will be twice as old as his son find the present age of his son (in years)
f + 15 = 2(s + 15)
f + 15 = 2s + 30
f = 2s + 30 - 15
f = 2s + 15
:
f=f, therefore
3s - 10 = 2s + 15
3s - 2s = 15 + 10
s = 25 yrs is the son's age
:
:
You can check this yourself, find the father's age and check both statements